Let 𝐼 = 010 𝑓(𝑥)𝑑x . You are asked to approximate the value of I using (i) uniform random 0 10 ∫ 𝑓(𝑥)𝑑𝑥 variable, (ii) Monte Carlo simulation and (iii) antithetic variates


a. Write down the procedures for the Monte Carlo simulation without using antithetic variates.


b. Write down the antithetic variable.


1
Expert's answer
2021-12-15T08:47:18-0500

(ii)

Monte Carlo approximation of the integral:


"\int f(x)dx=\frac{1}{n}\sum f(x_i)"


where xare are independent observations of X

if n = 10, then:


"\int^{10}_0 f(x)dx=\frac{1}{10}\sum f(x_i)"


iii)

 antithetic variates:


"\int f(x)dx=\frac{1}{n}\sum \frac{f(x_i)}{2}+\frac{1}{n}\sum \frac{f(1-x_i)}{2}"


if n = 10, then:


"\int^{10}_0 f(x)dx=\frac{1}{10}\sum \frac{f(x_i)}{2}+\frac{1}{10}\sum \frac{f(1-x_i)}{2}"


i)

using uniform random variable:


"\int^{b}_a f(x)dx=\frac{b-a}{n}\sum f(x_i)"


"\int^{10}_0 f(x)dx=\frac{10}{n}\sum f(x_i)"


a.

Monte Carlo simulation without using antithetic variates:

"\int f(x)dx=\frac{1}{n}\sum f(x_i)"


where xare are independent observations of X

if n = 10, then:


"\int^{10}_0 f(x)dx=\frac{1}{10}\sum f(x_i)"


b.

antithetic variable is "(1-x_i)"